# Exploiting Spatial Diversity in Earth-to-Satellite Quantum-Classical Communications

### v1

@@ I Introduction
Quantum communications via Low-Earth-Orbit (LEO) satellites have become a key enabler in extending, to a global scale, the unprecedented information security guaranteed by quantum physics. As an alternative to the Discrete Variable (DV) paradigm (which maps information to the discrete features of single photons), the Continuous-Variable (CV) paradigm (which encodes information onto the quadrature variables of optical fields) has been much touted as a promising candidate towards the practical implementation of quantum communications since CV quantum states can be readily generated and measured with high-speed and high-efficiency optical devices [1]. Within the CV paradigm, the recently proposed concept of Simultaneous Quantum-Classical communication (SQCC) [2, 3] represents a promising direction towards the simpler integration of global quantum and classical communication networks. Utilizing a form of quantum-classical signaling achieved by applying a small quantum modulation onto a large (classically-modulated) coherent state, SQCC allows quantum and classical communications to be implemented using the same communication infrastructure and the same optical pulse with minimum additional complexity. In addition to proof-of-principle experimental demonstrations via optical fiber (e.g., [4]), the feasibility of SQCC has been recently explored within the context of satellite-based quantum communications [5]. SQCC has the ability to embed multiple quantum protocols such as Quantum Key Distribution (QKD) [6], and Quantum Direct Communication (QDC) [7], within a wider class of classical communication protocols.
@@
Despite the predominant use of downlink channels, the construction of a global-scale quantum network also requires the use of uplink (i.e., Earth-to-satellite) channels. Depending on the specific use-case scenario, uplink channels are sometimes preferred due to their unique advantages, especially when the limitation of resources at a satellite and the flexibility of a ground-station system are taken into consideration (see discussions in, e.g., [14, 15, 16]). However, it is well-known that quantum-state transfer over an uplink channel is more challenging since the optical beam encounters atmospheric turbulence at the very start of its propagation with a beam width smaller than the outer scale of the turbulence (i.e. the upper cut off in the size of turbulent eddies) at the transmission point. The early-stage wavefront distortions imposed on the transmitted optical beam are exacerbated by the remaining transmission path to the satellite receiver, leading to significant beam wandering and beam broadening in the far field (see discussions in, e.g., [8, 9, 15, 17, 18, 16]). As a result, severe fading that significantly hinders the feasibility of FSO communications can be expected within an uplink channel. In extreme cases, the wandering beam may completely miss the small receiving aperture at the satellite, and the resulting “deep fades” can render FSO communications non-viable over an uplink channel.
@@
In FSO communications, despite the absence of multipath propagation, spatial diversity can still be introduced by adopting multiple transmitters and/or receivers since two subchannels spatially separated by a distance much larger than the atmospheric coherence length (usually on the order of centimeters) can be considered independent (see discussions in, e.g., [19, 20, 21]). Indeed, previous works have suggested the usefulness of utilizing spatial diversity in terrestrial (e.g., [22, 23, 24, 25, 26, 27]), satellite-to-Earth (e.g., [28, 29, 30, 31]), as well as Earth-to-satellite (e.g., [32, 33, 31]) FSO classical communications. The use of diversity has rarely been investigated within the context of quantum communications. The first step in this direction has been carried out recently in the work [34] which investigated the usefulness of spatial diversity in the discrimination of two coherent states in an analytical model of the fading channel. However, consideration of diversity usage for the purpose of quantum entanglement distribution through realistic Earth-to-satellite atmospheric fading channels remains unexplored – a critical missing element in the construction of a global quantum internet. Also missing is a detailed study of diversity-assisted coherent-state transfer over fading channels within the context of SQCC. In this work we deliver on both of these missing elements.
@@ II.1 System Configuration
As described in Fig. 1, in this work we consider an Earth-to-satellite (i.e., uplink) communication system with an $M\times M$ diversity scheme. In this system, $M$ coordinated ground stations, which can be located either within proximity on an observatory site or further from each other, send their corresponding quantum signals to a Low-Earth-Orbit (LEO) satellite employing an array of $M$ receiving apertures onboard. The transmitting ground stations and the receiving satellite apertures are matched in a one-to-one basis – specifically, transmitting ground station $j$ ($1\leq j\leq M$) is to send its signal beam to its designated receiving aperture (aperture $j$) at the satellite, and receiving aperture $j$ only expects the signal beam from its corresponding ground station (i.e., ground station $j$). We denote the coordinator of the $M$ ground station as Alice and the receiving satellite as Bob. Unless otherwise stated, we will assume that only the statistics, i.e., the Probability Density Function (PDF), of the transmissivity (as a function of channel direction) is known. This greatly simplifies the practical deployment of the communication system, and the required statistics can be achieved through some intensive a priori measurement scheme (in our calculations only Bob requires this information).
@@
Throughout this work, we assume that all the used subchannels are independent [^3] In order for the subchannels to be independent, the separations amongst the ground-station transmitters have to be much larger than the atmospheric coherence length. Since the atmospheric coherence length of the local atmosphere (i.e., several kilometers into the uplink atmospheric channel) at a ground-station is usually very small (our calculations using well-known equations in [8] give several centimeters), such separations can be easily achieved amongst the ground-station transmitters. since the transmitted beams from different ground stations evolve independently within their corresponding subchannels. As a result, the subchannel transmissivities $\{T_{j}\}$ can be modeled as independent random variables whose probability distributions are determined by the atmospheric optical turbulence. For simplicity, we will assume that all the subchannels are identically distributed, making $\{T_{j}\}$ independent and identically distributed (i.i.d.) random variables. Our channel modelling techniques are detailed in Appendix A with the fundamental theory of laser propagation within the Earth’s turbulent atmosphere presented in Appendix. 1.
@@ II.2 Operational Principles of Diversity-Assisted System
In an $M\times M$ diversity-assisted system utilizing $M$ subchannels, instead of directly transmitting a quantum state to Bob over a single channel, Alice equally splits the quantum state to be transmitted into $M$ parts using an array of beam splitters. She then transfers each of the $M$ parts to one of the $M$ involving ground stations, which, under Alice’s coordination, transmits its corresponding part through one of the $M$ subchannels. In this work, we assume the quantum state transfer between Alice and each of the other ground stations is perfect, i.e., achieved with no loss, noise, or decoherence. In this sense, “Alice” refers to all ground stations.
At the receiving satellite (Bob), the $M$ received signals (received by different receiving apertures) are coherently combined in the optical domain using a combining module consisting of multiple beam splitters. More details regarding the configuration and mathematical description of such a combining module are provided in Appendix B.
@@ III.1 Quantum State Evolution
We use the statistical-moment formalism (more details can be found in [38]) to describe the quantum-state evolution in entanglement distribution. In entanglement distribution without using diversity, Alice simply passes mode B of her TMSV state $\rho_{\text{A}{\text{B}}}$ to a ground station. The ground station then transmits mode B through an atmospheric channel (with channel transmissivity $T$ and excess noise $\epsilon$) to the satellite. As described in Sec. B of [38], we model the atmospheric channel using a beam splitter with the second input (representing the associated environmental mode) initialized in a thermal state. The Covariance Matrix (CM) of the total state (i.e., the TMSV state plus a single-mode thermal state) is given by
@@
where $V_{\text{s}}$ is the quadrature variance of the TMSV state, $\mathbb{1}_{2}=\operatorname{diag}(1,1)$, $\mathbb{Z}_{2}=\operatorname{diag}(1,-1)$, and $\oplus$ denotes a direct sum. After applying the beam splitter transformation $\mathbf{B}(\sqrt{T})$ corresponding to the atmospheric channel and tracing out the associated environmental mode (described in Sec. A of [38]), we can express the CM of the resulting two-mode state $\smash{\rho_{\text{A}{\text{B}}^{\prime}}^{(1)}}$ shared between Alice and Bob as
@@
After each part of mode B is sent though a subchannel to Bob, the resulting total quantum state (with an environmental mode associated with each subchannel taken into account) is a $\smash{(2M+1)}$-mode state. By tracing out all the $M$ environmental modes, we can obtain the $\smash{(M+1)}$-mode state after subchannel transmission. After taking into account Bob’s diversity combining (based on the configuration described in Appendix B) and tracing out all the unused modes, we can express the CM of the two-mode state $\rho_{\text{A}{\text{B}}^{\prime}}^{(M)}$ shared between Alice and Bob as
@@ III.2 Performance Metric
Intuitively, the performance of entanglement distribution is quantified by the amount of entanglement shared between Alice and Bob after Bob’s diversity combining, which is the amount of entanglement contained in the final two-mode state $\smash{\rho_{\text{A}{\text{B}}^{\prime},\text{avg}}^{(M)}}$. In this work, we adopt two of the most widely used entanglement measures, namely, the logarithmic negativity $E_{\text{LN}}$ and the Reverse Coherent Information (RCI) capacity $\mathcal{R}_{\text{C}}$, as our performance metrics for entanglement distribution. In particular, we evaluate these two entanglement measures using well-known methods assuming the two-mode state in question is a Gaussian state. Although the final two-mode state ${\rho}_{{\text{A}{\text{B}}^{\prime}},\text{avg}}^{(M)}$ shared between Alice and Bob is clearly a non-Gaussian mixed state, such a Gaussian approximation is still justified due to the extremality of Gaussian states – indeed, it is well-known that for every given CM, the entanglement is minimized by Gaussian states [40]. In other words, our result on the amount of distributed entanglement represents a lower bound on the amount [^4] Note that, the efficiently evaluation of the exact amount of entanglement contained in an arbitrary state is a highly non-trivial task. of entanglement contained in the final two-mode state ${\rho}_{{\text{A}{\text{B}}^{\prime}},\text{avg}}^{(M)}$ shared between Alice and Bob. Such a lower bound is commonly referred to as the Gaussian entanglement, and it quantifies the amount of entanglement that is useful for Gaussian quantum information protocols such as standard CV-QKD and the teleportation of Gaussian states (see discussions in, e.g., [41]).
@@ III.2.1 Logarithmic Negativity
III.2.2 RCI Capacity
As another useful entanglement measure that gives a lower bound of the distillable entanglement, the RCI (computed over Alice and Bob’s shared state $\rho_{\text{A}{\text{B}}}$) is given by
$I\left(\text{A}\!\shortleftarrow\!\text{B}\right)=H\left(\rho_{\text{A}}\right)-H\left(\rho_{\text{A}{\text{B}}}\right),$ (13)
where $H(\cdot)$ denotes von Neumann entropy [45]. In order to gain more fundamental insights into the use of diversity in entanglement distribution, in this work we evaluate the RCI capacity, which is obtained by maximizing RCI over all possible initial states [45, 46]. As shown by [46], for a single-mode Gaussian channel with transmissivity $T$, the RCI capacity is simply given by
$\mathcal{R}_{\text{C}}(T,\bar{n})=\max\left\{0,\log\left|\frac{1}{1-T}\right|-g(\bar{n})\right\},$ (14)
where $\bar{n}$ is the thermal number of the channel, and $g(x)=(x+1)\log_{2}(x+1)-x\log_{2}x$. In order to evaluate the RCI capacity within the single-mode framework, we recall from Eq. (8) that $M\times M$ diversity-assisted entanglement distribution over $M$ i.i.d. fluctuating subchannels without channel knowledge results in an effective non-fluctuating channel with effective transmissivity ${T_{\text{eff}}}$ and the effective thermal number $\smash{\bar{n}_{\text{eff}}=\frac{\frac{1}{M}\operatorname{Var}[\sqrt{T}](V_{\text{s}}-1)+\langle\epsilon\rangle-T_{\text{eff}}}{2(1-T_{\text{eff}})}-\frac{1}{2}}$ [47]. In other words, the RCI capacity achieved in our diversity-assisted entanglement distribution can be evaluated as $\mathcal{R}_{\text{C}}(T_{\text{eff}},\bar{n}_{\text{eff}})$.

### v2

@@ I Introduction
Quantum communications via Low-Earth-Orbit (LEO) satellites have become a key enabler in extending, to a global scale, the unprecedented information security guaranteed by quantum physics. As an alternative to the Discrete Variable (DV) paradigm (which maps information to the discrete features of single photons), the Continuous-Variable (CV) paradigm (which encodes information onto the quadrature variables of optical fields) has been much touted as a promising candidate towards the practical implementation of quantum communications since CV quantum states can be readily generated and measured with high-speed and high-efficiency optical devices [1]. Within the CV paradigm, the recently proposed concept of Simultaneous Quantum-Classical communication (SQCC) [2, 3] represents a promising direction towards the simpler integration of global quantum and classical communication networks. Utilizing a form of quantum-classical signaling achieved by applying a small quantum modulation onto a large (classically-modulated) coherent state, SQCC allows quantum communications and classical communications (including those supporting quantum communications) to be implemented using the same communication infrastructure (e.g., the same set of transmitter and receiver) and the same optical pulse with minimum additional complexity. In addition to proof-of-principle experimental demonstrations via optical fiber (e.g., [4]), the feasibility of SQCC has been recently explored within the context of satellite-based quantum communications [5]. SQCC has the ability to embed multiple quantum protocols such as Quantum Key Distribution (QKD) [6], and Quantum Direct Communication (QDC) [7], within a wider class of classical communication protocols.
@@
Despite the predominant use of downlink channels, the construction of a global-scale quantum network also requires the use of uplink (i.e., Earth-to-satellite) channels. Depending on the specific use-case scenario, uplink channels are sometimes preferred due to their unique advantages, especially when the limitation of resources at a satellite and the flexibility of a ground-station system are taken into consideration (see discussions in, e.g., [14, 15, 16]). Indeed, the uplink configuration avoids deploying high-quality CV quantum entanglement sources on satellites, which is, at the moment, still impractical from a system complexity point of view. From an operational point of view, the uplink configuration benefits from deploying quantum light sources at the ground stations, bringing great flexibility and allowing for future upgrades. In principle, a satellite-based system adopting the uplink configuration can support various quantum information protocols using different quantum light sources that can be readily diagnosed, maintained, reconfigured, or switched out at the ground stations. The satellite payload can also be more versatile (relative to one containing the quantum light source) since it may naturally be compatible with various quantum information protocols, including those requiring the satellite to play the simple role of a relay. Furthermore, although some well-known quantum information protocols can be implemented using solely downlink channels, the uplink configuration makes attacks directed at the receiver significantly more difficult in practice and is naturally required by many advanced quantum information protocols (a more detailed discussion regarding the usefulness of the uplink configuration in quantum communications is provided in Sec. S-III of [17]). However, it is well-known that quantum-state transfer over an uplink channel is more challenging since the optical beam encounters atmospheric turbulence at the very start of its propagation with a beam width smaller than the outer scale of the turbulence (i.e. the upper cut off in the size of turbulent eddies) at the transmission point. The early-stage wavefront distortions imposed on the transmitted optical beam are exacerbated by the remaining transmission path to the satellite receiver, leading to significant beam wandering and beam broadening in the far field (see discussions in, e.g., [8, 9, 15, 18, 19, 16]). As a result, severe fading that significantly hinders the feasibility of FSO communications can be expected within an uplink channel. In extreme cases, the wandering beam may completely miss the small receiving aperture at the satellite, and the resulting “deep fades” can render FSO communications non-viable over an uplink channel.
@@
In FSO communications, despite the absence of multipath propagation, spatial diversity can still be introduced by adopting multiple transmitters and/or receivers since two subchannels spatially separated by a distance much larger than the atmospheric coherence length (usually on the order of centimeters) can be considered independent (see discussions in, e.g., [20] for general FSO channels and [21, 22] for satellite-based FSO channels). Indeed, previous works have suggested the usefulness of utilizing spatial diversity in terrestrial (e.g., [23, 24, 25, 26, 27, 28]), satellite-to-Earth (e.g., [29, 30, 31, 32]), as well as Earth-to-satellite (e.g., [33, 34, 32]) FSO classical communications. The use of diversity has rarely been investigated within the context of quantum communications. The first step in this direction has been carried out recently in the work [35] which investigated the usefulness of spatial diversity in the discrimination of two small coherent states in an analytical model of the fading channel. However, it can be argued that the discrimination of two small coherent states in [35] has limited direct application to deployed CV-QKD. More importantly, consideration of diversity usage for the purpose of quantum entanglement distribution through realistic Earth-to-satellite atmospheric fading channels remains unexplored – a critical missing element in the construction of a global quantum internet. Also missing is a detailed study of diversity-assisted coherent-state transfer over fading channels within the context of SQCC. In this work we deliver on both of these missing elements.
@@
We should emphasize that although we study a combined quantum-classical communication system, in what is to follow the vast majority of our analysis concerns the quantum communications aspects. This is because the fragile nature of quantum signals, their inability to be copied, their quantum-only properties, and their complex evolution through a channel lead to a host of complexities absent in classical signalling.
@@ II.1 System Configuration
As described in Fig. 1, in this work we consider an Earth-to-satellite (i.e., uplink) communication system with an $M\times M$ diversity scheme. In this system, $M$ coordinated ground stations (located within proximity on an observatory site) send their corresponding quantum signals to a Low-Earth-Orbit (LEO) satellite employing an array of $M$ receiving apertures onboard. The transmitting ground stations and the receiving satellite apertures are matched in a one-to-one basis – specifically, transmitting ground station $j$ ($1\leq j\leq M$) is to send its signal beam to its designated receiving aperture (aperture $j$) at the satellite, and receiving aperture $j$ only expects the signal beam from its corresponding ground station (i.e., ground station $j$). We denote the coordinator of the $M$ ground station as Alice and the receiving satellite as Bob. Unless otherwise stated, we will assume that only the statistics, i.e., the Probability Density Function (PDF), of the transmissivity (as a function of channel direction) is known. This greatly simplifies the practical deployment of the communication system, and the required statistics can be achieved through some intensive a priori measurement scheme (in our calculations only Bob requires this information).
@@
Throughout this work, we assume that all the used subchannels are independent [^3] In order for the subchannels to be independent, the separations amongst the ground-station transmitters have to be much larger than the atmospheric coherence length. Since the atmospheric coherence length of the local atmosphere (i.e., several kilometers into the uplink atmospheric channel) at a ground-station is usually very small (our calculations using well-known equations in [8] give several centimeters), such separations can be easily achieved amongst the ground-station transmitters. since the transmitted beams from different ground stations evolve independently within their corresponding subchannels. As a result, the subchannel transmissivities $\{T_{j}\}$ can be modeled as independent random variables whose probability distributions are determined by the atmospheric optical turbulence. For simplicity, we will assume that all the subchannels are identically distributed, making $\{T_{j}\}$ independent and identically distributed (i.i.d.) random variables. Our channel modelling techniques are detailed in Appendix A with the fundamental theory of laser propagation within the Earth’s turbulent atmosphere presented in Appendix. 1. It should be noted that the statistical independence between the $M$ subchannels is a result of sufficient separation between the $M$ ground station transmitters. Since the coherence length at the satellite is of the order of $\sim\!1\,\text{km}$, it is practically infeasible to exploit diversity by employing a separation between the $M$ receiving apertures at a single satellite in an uplink configuration [21].
@@ II.2 Operational Principles of Diversity-Assisted System
In an $M\times M$ diversity-assisted system utilizing $M$ subchannels, instead of directly transmitting a quantum state to Bob over a single channel, Alice equally splits the quantum state to be transmitted into $M$ parts using an array of beam splitters. She then transfers each of the $M$ parts to one of the $M$ involving ground stations, which, under Alice’s coordination, transmits its corresponding part through one of the $M$ subchannels. In this work, we assume the quantum state transfer between Alice and each of the other ground stations is perfect, i.e., achieved with no loss, noise, or decoherence. In this sense, “Alice” refers to all ground stations. At the receiving satellite (Bob), the $M$ received signals (received by different receiving apertures) are coherently combined in the optical domain using a combining module consisting of multiple beam splitters. More details regarding the configuration and mathematical description of such a combining module are provided in Appendix C.
[…]
Here, we address our assumption of negligible beam-overlapping impact, which, in our scenario of interest, is not only well justified but also represents the worst-case scenario to some extent. In a practical implementation, we envision that high-precision tracking (demonstrated in the uplink quantum teleportation experiment [39]) is maintained between each pair of transmitting and receiving apertures. Since the beams from different subchannels have different angles/directions of arrival, from the perspective of each receiving aperture, all the subchannel beams except for one expected beam are tilted in their wavefronts. Since a wavefront tilt is converted into a spot displacement at the focal plane of the receiving optics, the portion of an unexpected beam that illuminates a receiving aperture (due to beam overlapping) cannot be effectively coupled into the system. As a result, beam overlapping at the satellite has a negligible impact on our system.
Non-negligible beam-overlapping impact arises when pairwise pointing and tracking cannot be maintained between every transmitting aperture and its corresponding receiving aperture. For example, if all the ground-station transmitting apertures adopt the same pointing angle (while the receiving apertures at the satellite point in different directions to achieve tracking), the assumption that different subchannels have different angles of arrival is valid only when the ground stations are separated by a large distance that may hinder the effectiveness of quantum-state transfer between Alice and the ground stations. The same is true when only one receiving aperture is used at the satellite, and all the transmitting apertures have to point to the same receiving aperture. However, rather than creating undesirable crosstalk noise, beam overlapping effectively couples more portions of signal into the system and allows for a potential performance enhancement since our scheme naturally requires all the subchannel signals to be combined at Bob to reproduce the very signal before Alice’s splitting. It should be noted that our diversity-assisted scheme, where different transmitters send exactly the same signals derived from an equal splitting of one single source, is fundamentally different from the scheme proposed in [40], where different transmitters send different signals independent of each other. In particular, although the scheme proposed in [40] leverages the full multi-input multi-output processing power (which is not always practically feasible in an uplink configuration), its performance is limited by the negative impact of beam overlapping (if not negligible) and its reliance on knowledge of instantaneous channel transmissivity for proper operation.
@@ III.1 Quantum State Evolution
We use the statistical-moment formalism (more details can be found in [17]) to describe the quantum-state evolution in entanglement distribution. In entanglement distribution without using diversity, Alice simply passes mode B of her TMSV state $\rho_{\text{A}{\text{B}}}$ to a ground station. The ground station then transmits mode B through an atmospheric channel (with channel transmissivity $T$ and excess noise $\epsilon$) to the satellite. As described in Appendix 2, we model the atmospheric channel using a beam splitter with the second input (representing the associated environmental mode) initialized in a thermal state. The Covariance Matrix (CM) of the total state (i.e., the TMSV state plus a single-mode thermal state) is given by
@@
where $V_{\text{s}}$ is the quadrature variance of the TMSV state, $\mathbb{1}_{2}=\operatorname{diag}(1,1)$, $\mathbb{Z}_{2}=\operatorname{diag}(1,-1)$, and $\oplus$ denotes a direct sum. After applying the beam splitter transformation $\mathbf{B}(\sqrt{T})$ corresponding to the atmospheric channel and tracing out the associated environmental mode (described in Appendix 1), we can express the CM of the resulting two-mode state $\smash{\rho_{\text{A}{\text{B}}^{\prime}}^{(1)}}$ shared between Alice and Bob as
@@
After each part of mode B is sent though a subchannel to Bob, the resulting total quantum state (with an environmental mode associated with each subchannel taken into account) is a $\smash{(2M+1)}$-mode state. By tracing out all the $M$ environmental modes, we can obtain the $\smash{(M+1)}$-mode state after subchannel transmission. After taking into account Bob’s diversity combining (based on the configuration described in Appendix C) and tracing out all the unused modes, we can express the CM of the two-mode state $\rho_{\text{A}{\text{B}}^{\prime}}^{(M)}$ shared between Alice and Bob as
@@ III.2 Performance Metric
Intuitively, the performance of entanglement distribution is quantified by the amount of entanglement shared between Alice and Bob after Bob’s diversity combining, which is the amount of entanglement contained in the final two-mode state $\smash{\rho_{\text{A}{\text{B}}^{\prime},\text{avg}}^{(M)}}$. In this work, we adopt two of the most widely used entanglement measures, namely, the logarithmic negativity $E_{\text{LN}}$ and the Reverse Coherent Information (RCI) $\mathcal{R}_{\text{C}}$, as our performance metrics for entanglement distribution. In particular, we evaluate these two entanglement measures using well-known methods assuming the two-mode state in question is a Gaussian state. Although the final two-mode state ${\rho}_{{\text{A}{\text{B}}^{\prime}},\text{avg}}^{(M)}$ shared between Alice and Bob is clearly a non-Gaussian mixed state, such a Gaussian approximation is still justified due to the extremality of Gaussian states – indeed, it is well-known that for every given CM, the entanglement is minimized by Gaussian states [42]. In other words, our result on the amount of distributed entanglement represents a lower bound on the amount [^4] Note that, the efficiently evaluation of the exact amount of entanglement contained in an arbitrary state is a highly non-trivial task. of entanglement contained in the final two-mode state ${\rho}_{{\text{A}{\text{B}}^{\prime}},\text{avg}}^{(M)}$ shared between Alice and Bob. Such a lower bound is commonly referred to as the Gaussian entanglement, and it quantifies the amount of entanglement that is useful for Gaussian quantum information protocols such as standard CV-QKD and the teleportation of Gaussian states (see discussions in, e.g., [43]).
@@ III.2.1 Logarithmic Negativity
III.2.2 RCI
As another useful entanglement measure that gives a lower bound of the distillable entanglement, the RCI (computed over Alice and Bob’s shared state $\rho_{\text{A}{\text{B}}}$) is defined as [47, 48]
$\mathcal{R}_{\text{C}}=H\left(\rho_{\text{A}}\right)-H\left(\rho_{\text{A}{\text{B}}}\right),$ (13)
where $\smash{\rho_{\text{A}}=\operatorname{tr}_{\text{B}}\left(\rho_{\text{A}{\text{B}}}\right)}$ and $H(\cdot)$ denotes the von Neumann entropy. In the spacial cases where $\rho_{\text{AB}}$ is a two-mode Gaussian state whose CM follows the general form given by Eq. (9), the RCI is simply given by
$\mathcal{R}_{\text{C}}=g(\nu^{(\text{A})})-\sum_{k}g\left(\nu_{k}^{(\text{AB})}\right),$ (14)
where $\smash{\nu^{(\text{A})}}$ denotes the symplectic eigenvalue of $\mathbf{V}_{\text{A}}$ (i.e., the CM of $\rho_{\text{A}}$), $\smash{\{\nu_{k}^{(\text{AB})}\}_{k=1}^{2}}$ denote the symplectic eigenvalues of $\mathbf{V}_{\text{AB}}$ (i.e., the CM of $\rho_{\text{AB}}$), and $\smash{g(x)=\frac{x+1}{2}\log_{2}(\frac{x+1}{2})-\frac{x-1}{2}\log_{2}(\frac{x-1}{2})}$. The symplectic eigenvalues $\smash{\{\nu_{k}\}_{k=1}^{N}}$ of a $2N$-by-$2N$ CM $\mathbf{V}$ can be readily computed as the modulus of the standard eigenvalues of the the matrix $\tilde{\mathbf{V}}=i\boldsymbol{\Omega}\mathbf{V}$, where $\boldsymbol{\Omega}$ is the symplectic form [49] given by Eq. (S3) of [17].
The distributed entanglement can be used for various quantum information protocols, the most well-known of which is QKD. Generally, in a CV-QKD protocol, after measurement and sifting, Alice and Bob perform parameter estimation, where they assume all the lost signals are acquired by Eve (who controls the channel) and determine the Holevo bound, other relevant parameters, and eventually, the secret key rate from their estimation of the CM in an EB picture [36]. In our scenario of interest, none of the estimated CM elements is a function of the instantaneous subchannel transmissivities or noises. Since we do not limit Eve’s capabilities in any way, all our analyses and findings are valid under the most powerful attacks allowed by quantum physics.
In this work, we focus on the performance of entanglement distribution rather than the performance of any specific quantum information protocol (that utilizes the distributed entanglement) since entanglement distribution is a more fundamental application that not only directly supports a variety of quantum information protocols (such as quantum teleportation and different QKD protocols) but also enables the establishment of entanglement between arbitrary network nodes (via entanglement swapping and distillation carried out by quantum repeaters [50]) over a distance much longer than that can be supported by a single satellite, underpinning the emergence of a true global-scale quantum internet.